OpenStudy (anonymous):

solution. can i please just have the answer. please. x2 + 10x + 25 = 8

OpenStudy (anonymous):

Just post all your homework here. Glad to do it for you.

OpenStudy (anonymous):

x2+35x=8 is wrong right.

OpenStudy (anonymous):

yes, they want you to solve for x

OpenStudy (anonymous):

yes well i can't even simplify from the original problem. i am trying to put it into the formula but I'm not doing it right.

OpenStudy (anonymous):

First you have to get all terms on one side and 0 on the other side

OpenStudy (anonymous):

subtract 25 from 8

OpenStudy (anonymous):

subtract 8 from 25

OpenStudy (anonymous):

okay so x^2+10x+8=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x^2+10x+25-8=0

OpenStudy (anonymous):

x^2+10x+17=0

OpenStudy (anonymous):

thats okay yeah my bad. i got that.

OpenStudy (anonymous):

so you have it in the form of ax^2+bx+c=0 what are a,b,c?

OpenStudy (anonymous):

a= x^2 b=10x c=0

OpenStudy (anonymous):

but the 17?

OpenStudy (anonymous):

just add it or...

OpenStudy (anonymous):

no, a, b, and c are coefficients, you don't include the variables

OpenStudy (anonymous):

a=1, b=10, c=17

OpenStudy (anonymous):

alright. i got that. thanks for going through this with my btw...

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

o.o alright can you step me through that i don't know what the sign is after the b which is -10 right?

OpenStudy (anonymous):

thats a plus or minus, means there are 2 solutions, similar to this:\[x=\frac{-b + \sqrt{b^2-4ac}}{2a}, \frac{-b - \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

-b is -10

OpenStudy (anonymous):

ah. alright that makes sense.

OpenStudy (anonymous):

\[x=\frac{-10 + \sqrt{10^2-4(1)(17)}}{2(1)}\] and \[x=\frac{-10 - \sqrt{10^2-4(1)(17)}}{2(1)}\]

OpenStudy (anonymous):

alright so all i do is "do it" right?

OpenStudy (anonymous):

yea, simplify it

OpenStudy (anonymous):

alright. so 10 squared is 100 - 4 = 96 multiplied by 17? i can't be doing this right

OpenStudy (anonymous):

PEMDAS, multiplication and division come before addition and subtraction

OpenStudy (anonymous):

okay so i square the 10 then times 17 by 1 and then by four?

OpenStudy (anonymous):

100-[4*1*17]

OpenStudy (anonymous):

32

OpenStudy (anonymous):

\[x=\frac{-10 \pm \sqrt{32}}{2}\] simplifiy more, the radical can be simplified

OpenStudy (anonymous):

the sqr of 32 is 5 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:I six if i round?

OpenStudy (anonymous):

\[\sqrt{32}=\sqrt{16*2}=\sqrt{16}*\sqrt{2}=4\sqrt{2}\]

OpenStudy (anonymous):

ah... okay and then i ...

OpenStudy (anonymous):

simplify more

OpenStudy (anonymous):

can i add at this point because i can't exactly divide?

OpenStudy (anonymous):

\[x=\frac{-10 \pm 4\sqrt{2}}{2}=-5 \pm 2\sqrt{2}\]

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

thank you for stepping me through that dockworker

OpenStudy (anonymous):

you're welcome